What mass of CuO can be formed from the reaction of 8.35g of CuS and 7.20g of O2


1
Expert's answer
2021-03-08T06:06:05-0500

Solution:

The balanced chemical equation:

2CuS + 3O2 → 2CuO + 2SO2

According to the equation above: n(CuS) = 2×n(O2)/3 = n(CuO)


Let's start by converting the masses of CuS and O2​ to moles using their molar masses:

8.35 g CuS × (1 mol CuS / 95.611 g) = 0.08733 mol CuS

7.20 g O2​ × (1 mol O2​ / 31.9988 g) = 0.2250 mol O2


Choose one reactant and determine how many moles of the other reactant are necessary to completely react with it. Let's choose CuS:

n(O2) = 3 × n(CuS) / 2 = (3 × 0.08733 mol) / 2 = 0.130995 mol​


The calculation above means that we need 0.130995 mol of O2 to completely react with CuS.

We have 0.2250 mol of O2 and therefore more than enough oxygen.

Thus oxygen (O2) is in excess and copper monosulfide (CuS) must be the limiting reagent.


Thus: n(CuO) = n(CuS) = 0.08733 mol


Moles of CuO = Mass of CuO / Molar mass of CuO

Mass of CuO = Moles of CuO × Molar mass of CuO

The molar mass of CuO is 79.545 g mol-1.

Hence,

Mass of CuO = 0.08733 mol × 79.545 g mol-1 = 6.9466 g = 6.95 g

Mass of CuO = 6.95 g


Answer: 6.95 grams of CuO can be formed.